How many liters of an 80% glycol solution must be added to 6 liters of a 25% glycol solution to obtain a 55% solution?
Answer by richwmiller(17219)

0.8*x+0.25*6=0.55(6+x)
0.8*x+1.5=3.3+0.55x
0.8x-0.55x=3.3-1.5
0.25x=1.8
x=7.2 liter of 80% is added to 6 liter of 25% to get 13.2 liter at 55%
check
0.8*7.2+0.25*6=0.55(6+7.2)
5.76+1.5=0.55(13.2)
7.26=7.26
ok